# Model Predictive Control-Based Optimal Energy Management of Autonomous Electric Vehicles Under Cold Temperatures

### v1

@@ Abstract
keywords:
Energy Management, Cabin Heating, HVAC, Batteries, Autonomous Electric Vehicle, Model Predictive Control
journal: Applied Energy
{highlights}
MPC-based optimal power split between vehicle propulsion, HVAC, and battery thermal management is designed
Heat pump mode of operation of HVAC for Cabin heating is used for the HVAC architecture
Case studies for low SoC and cold ambient temperatures were conducted
Open-source HVAC properties from CoolProp have been used
Open Street Maps and OpenTopology were used for the real-world road grade emulation
1 Introduction
1.1 Motivation
@@ 1.2 Literature Review
1.3 Contribution and Summary of the Paper
@@
2 Energy Management Problem Under Cold Temperature
@@
3 Mathematical Modeling of Autonomous Electric Vehicle Subsystems
@@
3.1 Vehicle Propulsion Subsystem
@@
where $m_{vehicle}$ is the vehicle mass in $[kg]$, $F_{propulsion}$ is the propulsion force that is powered by the battery in $[N]$, $\mu$ is the dimensionless road’s friction coefficient, $\rho_{air}$ is the air density in $[kg/m^{3}]$, $\theta$ is the road grade in $[rads]$, $A_{vehicle}$ is the frontal surface area of the vehicle that is exposed to the air drag in $[m^{2}]$ and $C_{d}$ is the dimensionless coefficient of drag force. $x_{dist}$ is the distance in $[m]$ traveled by the vehicle in $dt$ timestep. Furthermore, the energy spent in providing the propulsion force $F_{propulsion}$ is given by:
[…]
3.2 Cabin Thermal Subsystem
@@
3.3 HVAC Subsystem
In this subsection, we describe the HVAC subsystem. Figure 2 depicts the architecture of the refrigeration loop, coolant loop and the battery heater, which has been adopted after referring to [24]. Under cold temperatures, the cabin and battery temperature requests are intended to provide heat to the cabin as well as the battery. This means that the HVAC must operate in the heating mode, which is also commonly referred to as the heat pump mode of operation. It can be seen in Fig. 2 that the refrigerant enters the compressor at low temperature and low pressure (State 1). It is assumed that the properties of State 1 are fully known. Depending upon the requested temperature setting from the user/controller, the required compressor’s work input is estimated, thereby State 2 is predicted, and the refrigerant is made to attain State 2. After this, the refrigerant goes into the condenser, where there is a heat exchange between the refrigerant and the air. The air takes heat from the refrigerant and cools it until the refrigerant becomes a slightly sub-cooled liquid at State 3. Next, it passes through an expansion valve where the pressure and temperature drop, and the refrigerant reaches State 4. Finally, it passes through an evaporator where it takes heat and reaches State 1 again. Typically, in the HVAC systems, based on the ambient temperature and the target air temperature, States 1 and 3 are estimated. Usually, the temperature of the refrigerant is about $5C$ less than the ambient temperature at State 1, while at State 3, the refrigerant is around $5-10C$ more than the target cabin temperature.
@@ 3.4 Battery Subsystem
where $R_{0}$ is the series resistance, $R_{1}$ is the resistance parallel to the capacitance both in $[\Omega]$; $C_{1}$ is the capacitance in $[F]$; $V_{c}$ is the voltage across the capacitor, $V_{Terminal}$ is the terminal voltage, and $OCV$ in the open circuit voltage – all of them in $[V]$; $Q$ is the battery capacity in $[A\-s]$ and $SoC$ is the state of charge (dimension-less quantity in $[0,1]$). $I_{demand}$ is the total current demand in $A$ from all the power requirements at the cell level.
@@
where $T_{Batt}$ and $T_{ambient}$ are battery cell and ambient temperatures in [$K$], $Q_{gen}$ is the heat generation due to vehicle propulsion-related battery current, $h_{Batt}$ is the battery convective heat transfer coefficient in [$Wm^{2}K^{-1}$], $A_{Batt}$ is surface area of the cell in [$m^{2}$], $E_{heater}=I_{heater}^{2}R_{heater}$ is the power spent to heat the battery for charging, with $I_{heater}$ being the current to the heater in [$A$] and $R_{heater}$ being the heater’s resistance in [$\Omega$], $m_{Batt}$ is the battery cell’s mass in $[kg]$ and $C_{p_{Batt}}$ is its heat capacity in [$Jkg^{-1}K^{-1}$].
3.5 Lower-Level Controllers
@@
4 Model Predictive Control for Optimal Energy Management
@@
The MPC formulation is run in an outer loop System-Level Controller, which is updated every $dt$ seconds to give out its output at the $k^{th}$ timestep, with a moving horizon prediction for the next $N$ timesteps. Out of these $N$ steps, only the first timestep is used as the input to the AEV lower-level controllers, discarding all the other $N-1$ predictions. This process continues till the very end of the vehicle operation. The perception system of the AEV – including cameras, Global Positioning System (GPS), Inertial Measurement Units (IMUs), outside temperature sensor – estimates the state of the environment for the next $N$ timesteps and provide this information to the MPC. The MPC then further use this information to minimize a cost function for the next $N$ timesteps. The underlying cost function that is optimized for the entire moving horizon of $N$ timesteps is as follows:
@@
where $y$ is the MPC output vector, $f_{1}(.)$, $f_{2}(.)$ and $f_{3}(.)$ are the linear functions described by (8), (10), and (1). The inequality constraints of the system states are given by (4), and the constraints on the MPC outputs are given by (4).
@@
5 Simulation Results and Discussion
@@
  • Battery cell is simulated considering a lithium-ion battery cell with 4.2V to 2.5V operating voltages with a rated capacity of 4000 mAh. The cell’s open circuit voltage ($OCV$) has been experimentally investigated by charging and discharging the battery cell at low C-rates. OCV is used in BMS to calibrate SOC estimations and make informed decisions to degrade the battery at a slower rate. Subsequently, the equivalent circuit parameters $Q$, $R_{0}$, $R_{1}$ and $C_{1}$ are calculated using the Reference Performance Test (RPT) and the identified parameter values are as follows: (1) $Q=1.4322\times 10^{4}[A-s]$; (2) $R_{0}=1.3513\times 10^{-2}[\Omega]$; (3) $R_{1}=1.028\times 10^{-2}[\Omega]$; (4) $C_{1}=5.2584\times 10^{3}[F]$ [31]. A battery pack is then modeled using this cell using 100S56P pack architecture, bringing the battery pack’s nominal voltage to around 400V. This is the voltage that powers the motors and all the power electronics, some of which might require a voltage step-down. However, all the sub-systems of current interest operate at 400V [32]. Furthermore, battery cell thermal dynamics is simulated using the properties listed in Table 1.
@@
  • Cabin Thermal system was modeled using the energy thermal balance as described in (3), where the ambient cold temperatures create the thermal gradient leading to heat transfer from the cabin to the environment through the vehicle body surfaces. The thermal parameters are listed in Table 2.
  • Vehicle Propulsion system was simulated using parameters taken from multiple sources including [33, 34], and the values of the associated properties are listed in Table 2.
Table 1: Battery Cell thermal parameters
@@
$h_{Batt}$ Car Heat transfer coeff. $10$ $[W/(m^{2}K)]$
$A_{Batt}$ Batt. surface area $5.3\times 10^{-3}$ $[m^{2}]$
$R_{Heater}$ Heater coeff. $4$ $[J/Ks]$
Table 2: Vehicle model parameters
@@
$A_{Frontal}$ Vehicle’s frontal area $2.2$ $[m^{2}]$
@@
$V_{cabin}$ Volume of cabin $2.5$ $[m^{3}]$
$\rho_{air}$ Air density $1.2$ $[kg/m^{3}]$
$A_{Car}$ Losses surface area $7$ $[m^{2}]$
$h_{car}$ Car Heat transfer coeff. $3.5$ $[W/(m^{2}K)]$
@@
5.1 Case Study 1
In this study, we simulate the environment using the ambient temperature profile and road grade as shown in Fig. 4. These environmental characteristics are chosen to capture some key outcomes of the proposed control scheme. Essentially, the ambient temperature oscillates between 268 $K$ and 278 $K$ during vehicle operation, which is assumed to last 600 seconds. Along this vehicular trajectory, it is also assumed that the road grade oscillates between $-20\%$ and $+20\%$. Furthermore, at the beginning of the simulation, the vehicle is moving at a constant velocity of $65MPH$, and the battery $SoC$ is $0.2$. The battery and cabin temperatures are at $293K$ and $278K$. The nearest charging station is located $10.8$ miles away, and the battery needs to be between $313.15K$ and $315.15K$ temperatures to perform efficient charging once it reaches the charging station.
@@ 5.1 Case Study 1
In Fig. 7, the velocity response dictated by the MPC controller is shown. The proposed controller uses the look-ahead information from the AEV’s sensors to estimate the road grade and ambient temperature. This is used to optimize propulsion, thereby leading to a time-varying speed profile as opposed to a constant speed profile. Since the vehicle is going downhill initially from 00 seconds to $80$ seconds, the vehicle speed increase is seen to be from $65MPH$ to around $68.5MPH$. But since for those $80$ seconds the power requirement is from all the three subsystems, the vehicle speed does not go beyond $68.5MPH$ as the energy gained in downhill is split across all the subsystems. During the vehicle’s uphill trajectory from $80$ seconds to $180$ seconds, the vehicle speed decreases until it reaches the minimum allowed speed on the freeway and subsequently maintains this speed. Unlike the first $80$ seconds, for the next wave of downhill in the road grade from $180$ seconds to $280$ seconds, the major energy consumption requirement is from the propulsion as the cabin and battery temperatures have already reached close to the targeted values. This is reflected in the vehicle speed as it reaches a higher value of around $70MPH$. However, due to the loss of energy due to inefficiency, in energy regeneration during the downhills, the momentum gained during uphills is lost, making the main priority to maintain the vehicle speed at the minimum allowed speed during all the following waves of the road, as energy regenerated by the vehicle during downhills is no longer sufficient to increase the vehicle speed to significant speeds. Therefore, in all the following waves from $400s$ to $600s$, the vehicle speed almost remains at the lowest allowed speed.
@@ 5.2 Case Study 2
Essentially, the ambient temperature oscillates between 256.65 $K$ and 259.65 $K$ during vehicle operation, which is assumed to last 600 seconds. Along this vehicular trajectory, it is also assumed that the road grade oscillates between an instantaneous grade of $-62.85\%$ and $+39.23\%$. Furthermore, at the beginning of the simulation, the vehicle is moving at a constant velocity of $65MPH$, and the battery $SoC$ is $0.2$. The battery and cabin temperatures are at $293K$ and $278K$. Before reaching the end destination as shown in Fig. 8, the battery needs to be between $313.15K$ and $315.15K$ temperatures to perform efficient charging once it reaches the charging station.
@@
6 Conclusion and Future Work

### v3

@@ Abstract
I Introduction
I-A Motivation
@@ I-B Literature Review
I-C Contribution and Summary of the Paper
@@
II Energy Management Problem Under Cold Temperature
@@
III Mathematical Modeling of Autonomous Electric Vehicle Subsystems
@@
III-A Vehicle Propulsion Subsystem
@@
where $m_{vehicle}$ is the vehicle mass in $[kg]$, $F_{propulsion}$ is the propulsion force that is powered by the battery in $[N]$, $\mu$ is the dimensionless road’s friction coefficient, $\rho_{air}$ is the air density in $kg/m^{3}]$, $\theta$ is the road grade in $[rads]$, $A_{vehicle}$ is the frontal surface area of the vehicle that is exposed to the air drag in $m^{2}]$ and $C_{d}$ is the dimensionless coefficient of drag force. $x_{dist}$ is the distance in $[m]$ traveled by the vehicle in $dt$ timestep. Furthermore, the energy spent in providing the propulsion force $F_{propulsion}$ is given by:
$E_{propulsion}=\int_{0}^{t}F_{propulsion}v_{vehicle}dt.$ (2)
III-B Cabin Thermal Subsystem
@@
III-C HVAC Subsystem
In this subsection, we describe the HVAC subsystem. Figure 2 depicts the architecture of the refrigeration loop, coolant loop and the battery heater, which has been adopted after referring to [24]. Under cold temperatures, the cabin and battery temperature requests are intended to provide heat to the cabin as well as the battery. This means that the HVAC must operate in the heating mode, which is also commonly referred to as the heat pump mode of operation. It can be seen in Fig. 2 that the refrigerant enters the compressor at low temperature and low pressure (State 1). It is assumed that the properties of State 1 are fully known. Depending upon the requested temperature setting from the user/controller, the required compressor’s work input is estimated, thereby State 2 is predicted, and the refrigerant is made to attain State 2. After this, the refrigerant goes into the condenser, where there is a heat exchange between the refrigerant and the air. The air takes heat from the refrigerant and cools it until the refrigerant becomes a slightly sub-cooled liquid at State 3. Next, it passes through an expansion valve where the pressure and temperature drop, and the refrigerant reaches State 4. Finally, it passes through an evaporator where it takes heat and reaches State 1 again. Typically, in the HVAC systems, based on the ambient temperature and the target air temperature, States 1 and 3 are estimated. Usually, the temperature of the refrigerant is about $5C$ less than the ambient temperature at State 1, while at State 3, the refrigerant is around 510C5-10C more than the target cabin temperature.
@@ III-D Battery Subsystem
where $R_{0}$ is the series resistance, $R_{1}$ is the resistance parallel to the capacitance both in $\Omega]$; $C_{1}$ is the capacitance in $[F]$; $V_{c}$ is the voltage across the capacitor, $V_{Terminal}$ is the terminal voltage, and $OCV$ in the open circuit voltage – all of them in $[V]$; $Q$ is the battery capacity in [A][A]-$[s]$ and $SoC$ is the state of charge (dimension-less quantity in $[0,1]$). $I_{demand}$ is the total current demand in $A$ from all the power requirements at the cell level.
@@
where $T_{Batt}$ and $T_{ambient}$ are battery cell and ambient temperatures in [$K$], $Q_{gen}$ is the heat generation due to vehicle propulsion-related battery current, $h_{Batt}$ is the battery convective heat transfer coefficient in $Wm^{2}K^{-1}$], $A_{Batt}$ is surface area of the cell in $m^{2}$], $E_{heater}=I_{heater}^{2}R_{heater}$ is the power spent to heat the battery for charging, with $I_{heater}$ being the current to the heater in [$A$] and $R_{heater}$ being the heater’s resistance in $\Omega$], $m_{Batt}$ is the battery cell’s mass in $[kg]$ and $C_{p_{Batt}}$ is its heat capacity in $Jkg^{-1}K^{-1}$].
III-E Lower-Level Controllers
@@
IV Model Predictive Control for Optimal Energy Management
@@
The MPC formulation is run in an outer loop System-Level Controller, which is updated every $dt$ seconds to give out its output at the $k^{th}$ timestep, with a moving horizon prediction for the next $N$ timesteps. Out of these $N$ steps, only the first timestep is used as the input to the AEV lower-level controllers, discarding all the other N1N-1 predictions. This process continues till the very end of the vehicle operation. The perception system of the AEV – including cameras, Global Positioning System (GPS), Inertial Measurement Units (IMUs), outside temperature sensor – estimates the state of the environment for the next $N$ timesteps and provide this information to the MPC. The MPC then further use this information to minimize a cost function for the next $N$ timesteps. The underlying cost function that is optimized for the entire moving horizon of $N$ timesteps is as follows:
@@
where $y$ is the MPC output vector, $f_{1}(.)$, $f_{2}(.)$ and $f_{3}(.)$ are the linear functions described by (8), (10), and (1). The inequality constraints of the system states are given by (IV), and the constraints on the MPC outputs are given by (IV).
@@
V Simulation Results and Discussion
@@
  • Battery cell is simulated considering a lithium-ion battery cell with 4.2V to 2.5V operating voltages with a rated capacity of 4000 mAh. The cell’s open circuit voltage ($OCV$) has been experimentally investigated by charging and discharging the battery cell at low C-rates. OCV is used in BMS to calibrate SOC estimations and make informed decisions to degrade the battery at a slower rate. Subsequently, the equivalent circuit parameters $Q$, $R_{0}$, $R_{1}$ and $C_{1}$ are calculated using the Reference Performance Test (RPT) and the identified parameter values are as follows: (1) $Q=1.4322\times 10^{4}[A-s]$; (2) $R_{0}=1.3513\times 10^{-2}[\Omega]$; (3) $R_{1}=1.028\times 10^{-2}[\Omega]$; (4) $C_{1}=5.2584\times 10^{3}[F]$ [31]. A battery pack is then modeled using this cell using 100S56P pack architecture, bringing the battery pack’s nominal voltage to around 400V. This is the voltage that powers the motors and all the power electronics, some of which might require a voltage step-down. However, all the sub-systems of current interest operate at 400V [32]. Furthermore, battery cell thermal dynamics is simulated using the properties listed in Table I.
@@
  • Cabin Thermal system was modeled using the energy thermal balance as described in (3), where the ambient cold temperatures create the thermal gradient leading to heat transfer from the cabin to the environment through the vehicle body surfaces. The thermal parameters are listed in Table II.
  • Vehicle Propulsion system was simulated using parameters taken from multiple sources including [33, 34], and the values of the associated properties are listed in Table II.
TABLE I: Battery Cell thermal parameters
@@
$h_{Batt}$ Car Heat transfer coeff. $10$ $W/(m^{2}K)]$
$A_{Batt}$ Batt. surface area $5.3\times 10^{-3}$ $m^{2}]$
$R_{Heater}$ Heater coeff. 44 $[J/Ks]$
TABLE II: Vehicle model parameters
@@
$A_{Frontal}$ Vehicle’s frontal area $2.2$ $m^{2}]$
@@
$V_{cabin}$ Volume of cabin $2.5$ $m^{3}]$
$\rho_{air}$ Air density $1.2$ $kg/m^{3}]$
$A_{Car}$ Losses surface area 77 $m^{2}]$
$h_{car}$ Car Heat transfer coeff. $3.5$ $W/(m^{2}K)]$
@@
V-A Case Study 1
In this study, we simulate the environment using the ambient temperature profile and road grade as shown in Fig. 4. These environmental characteristics are chosen to capture some key outcomes of the proposed control scheme. Essentially, the ambient temperature oscillates between 268 $K$ and 278 $K$ during vehicle operation, which is assumed to last 600 seconds. Along this vehicular trajectory, it is also assumed that the road grade oscillates between $20\%$ and $20\%$. Furthermore, at the beginning of the simulation, the vehicle is moving at a constant velocity of $65MPH$, and the battery $SoC$ is $0.2$. The battery and cabin temperatures are at $293K$ and $278K$. The nearest charging station is located $10.8$ miles away, and the battery needs to be between $313.15K$ and $315.15K$ temperatures to perform efficient charging once it reaches the charging station.
@@ V-A Case Study 1
In Fig. 7, the velocity response dictated by the MPC controller is shown. The proposed controller uses the look-ahead information from the AEV’s sensors to estimate the road grade and ambient temperature. This is used to optimize propulsion, thereby leading to a time-varying speed profile as opposed to a constant speed profile. Since the vehicle is going downhill initially from 0 seconds to $80$ seconds, the vehicle speed increase is seen to be from $65MPH$ to around $68.5MPH$. But since for those $80$ seconds the power requirement is from all the three subsystems, the vehicle speed does not go beyond $68.5MPH$ as the energy gained in downhill is split across all the subsystems. During the vehicle’s uphill trajectory from $80$ seconds to $180$ seconds, the vehicle speed decreases until it reaches the minimum allowed speed on the freeway and subsequently maintains this speed. Unlike the first $80$ seconds, for the next wave of downhill in the road grade from $180$ seconds to $280$ seconds, the major energy consumption requirement is from the propulsion as the cabin and battery temperatures have already reached close to the targeted values. This is reflected in the vehicle speed as it reaches a higher value of around $70MPH$. However, due to the loss of energy due to inefficiency, in energy regeneration during the downhills, the momentum gained during uphills is lost, making the main priority to maintain the vehicle speed at the minimum allowed speed during all the following waves of the road, as energy regenerated by the vehicle during downhills is no longer sufficient to increase the vehicle speed to significant speeds. Therefore, in all the following waves from $400s$ to $600s$, the vehicle speed almost remains at the lowest allowed speed.
@@ V-B Case Study 2
Essentially, the ambient temperature oscillates between 256.65 $K$ and 259.65 $K$ during vehicle operation, which is assumed to last 600 seconds. Along this vehicular trajectory, it is also assumed that the road grade oscillates between an instantaneous grade of $62.85\%$ and $39.23\%$. Furthermore, at the beginning of the simulation, the vehicle is moving at a constant velocity of $65MPH$, and the battery $SoC$ is $0.2$. The battery and cabin temperatures are at $293K$ and $278K$. Before reaching the end destination as shown in Fig. 8, the battery needs to be between $313.15K$ and $315.15K$ temperatures to perform efficient charging once it reaches the charging station.
@@
VI Conclusion and Future Work